OpenStudy (anonymous):

application of differentiation??? f(x)=2x^3-3x^2-12x+1; [-2,3]

OpenStudy (anonymous):

complete the question...

OpenStudy (anonymous):

Find the absolute mac and absolute min values of f on the given interval

OpenStudy (amistre64):

y' = 6x^2 -6x -12 tells us the max and min of the graph on all points. x^2 -x -2 = 0 (x-2)(x+1) = 0 when x=2 and x=-1 these are between [-2,3] y'' = 12x -6 tells is curved up(+) curved down(-) 12(2) -6 = 24-6 = positive, its a min (curved up) 12(-1) -6 = -12-6 = negative its a max (curved down) but I wonder if they also want to include the end point of the graph. cause they would most likely be max @ 3 and min@-2 solve for f(-2), f(-1), f(2), and f(3) to see which value is biggest :)

OpenStudy (anonymous):

omg m hero.....do you do tutorials...i started the problem and got exactly what you got but i get confused at time and need major help with my math skills?

OpenStudy (amistre64):

tutorials?...not that I am aware of :) I actually am self taught in calc. gonna take it over the summer semester in college tho so I have a foothold on it already.. If you have any questions, I might be able to help with them :)

OpenStudy (anonymous):

if your good in calculus that means your pretty good in college algebra and trig CORRECT????

OpenStudy (amistre64):

nothin but As :)

OpenStudy (anonymous):

Jealous so will you help me occasionally as much as possible...and if m grades improve ill pay you. IT IS THAT CRUCIAL :( sad to say

OpenStudy (amistre64):

I dont know how pay would work, but I will help you when I can.

OpenStudy (anonymous):

do you have an email address i can reach you at? and back to the problem did you take the derivative twice

OpenStudy (amistre64):

I usually dont take the serivative twice on a cubic equation because the graph looks like "N". but yeah, i figured I needed the practice so I took the derivative twice to see how it bent at those points. The first derivative gives you the bends (lol) and the second derivative tells you if it is bending up or down. If it is (+) it bends up like a bowl sitting on the table. If it is (-) it bends down like an upside down bowl.

OpenStudy (amistre64):

use my screenname and add "@gmail.com" for my email addy. I only have internet at school so I would be limited to when I could respond.

OpenStudy (anonymous):

thanks. working on problems now maybe i can do a couple problems and email them to you so you can check over them or places i get stuck and expect and email from dorothyclayton@yahoo.com thanks again you are greatly appreciated

OpenStudy (amistre64):

your welcome :)